OpenStudy (anonymous):

derivation of (xsinx)/(1+tanx)?please help me!

OpenStudy (anonymous):

Low(DerivativeHIGH)-High(DerivativeLOW)/Low*Low

OpenStudy (anonymous):

sorry!can u explain by step for me?

OpenStudy (amistre64):

that WAS the explanation :)

OpenStudy (amistre64):

its the application of the quotient rule

OpenStudy (anonymous):

that is the form!

OpenStudy (amistre64):

and the product rule since the top is a product

OpenStudy (amistre64):

\[\frac{d}{dx}\ \frac{x\ sinx}{1+tanx}\] \[ \frac{(1+tanx)(x'\ sinx+x\ sin'x)-(1+tanx)'(x\ sinx)}{(1+tanx)^2}\]

OpenStudy (amistre64):

play out the 's and your done

OpenStudy (anonymous):

ok ! thx so much!

OpenStudy (amistre64):

yep